# II-MMR: Identifying and Improving Multi-modal Multi-hop Reasoning in Visual Question Answering

### v1

@@ Abstract
Visual Question Answering (VQA) often involves diverse reasoning scenarios across Vision and Language (V&L). Most prior VQA studies, however, have merely focused on assessing the model’s overall accuracy without evaluating it on different reasoning cases. Furthermore, some recent works observe that conventional Chain-of-Thought (CoT) prompting fails to generate effective reasoning for VQA, especially for complex scenarios requiring multi-hop reasoning. In this paper, we propose II-MMR, a novel idea to identify and improve multi-modal multi-hop reasoning in VQA. In specific, II-MMR takes a VQA question with an image and finds a reasoning path to reach its answer using two novel language promptings: (i) answer prediction-guided CoT prompt, or (ii) knowledge triplet-guided prompt. II-MMR then analyzes this path to identify different reasoning cases in current VQA benchmarks by estimating how many hops and what types (i.e., visual or beyond-visual) of reasoning are required to answer the question. On popular benchmarks including GQA and A-OKVQA, II-MMR observes that most of their VQA questions are easy to answer, simply demanding “single-hop” reasoning, whereas only a few questions require “multi-hop” reasoning. Moreover, while the recent V&L model struggles with such complex multi-hop reasoning questions even using the traditional CoT method, II-MMR shows its effectiveness across all reasoning cases in both zero-shot and fine-tuning settings.
@@
[…]
@@ 1 Introduction
Effectiveness of II-MMR. II-MMR analyzes this reasoning path to identify different reasoning cases in current VQA benchmarks by measuring the number of reasoning steps and the types of reasoning (i.e., visual or beyond-visual) required to answer each question (Figure 1). During our prompting process, the intermediate reasoning steps and the alignment of question keywords with visual objects determine the number and types of reasoning.
@@ 3 Experimental Setup
Training details. We follow the official configurations and implementations of BLIP-2111https://github.com/salesforce/LAVIS/tree/main. Specifically, we select the largest BLIP-2 model, which its vision encoder is ViT-g Fang et al. (2023) and its LLM is FlanT5XXL Chung et al. (2022). During fine-tuning, the parameters of BLIP-2 are optimized with the language modeling loss on the downstream VQA training data with a batch size of 16 and a learning rate of 1e-5 for ten epochs. We utilize eight A100 48GB GPUs for both training and inference. See the appendix for more details.
@@ 4.1 Analysis of reasoning in VQA benchmarks
Table 6: Effectiveness of II-MMR over different reasoning cases in zero-shot VQA. Our II-MMR outperforms the traditional CoT on every reasoning case. Moreover, II-MMR notably improves over the baseline (BLIP-2) in all reasoning cases on both benchmarks, suggesting the benefit of our answer reasoning path in finding correct answers.
@@ 4.4 Benefit of II-MMR in fine-tuning stage
4.5 Expanding questions with more reasoning
As depicted in Table 2, the number of complex reasoning questions (i.e., 2-hop reasoning) is marginal in the GQA test-dev set, comprising only 3.73%. We thus conduct an ablation study: increasing the number of hops for each original question using a large-scale knowledge-base (e.g., Wikipedia) and evaluating the model performance on these newly expanded questions. At a high level, we first extract keywords from the original question and use them as queries to retrieve their relevant information from Wikipedia. We then input this extra information (with the original question) into the LLM to increase the reasoning complexity of the original question. See the appendix for more details.
@@ 4.6 Qualitative Results
Figure 5 shows some qualitative examples provided by our II-MMRApCoT. The generated rationale is highly relevant to the correct answer, leading the model to make the correct prediction. Moreover, our rationales entail knowledge beyond images, such as commonsense (e.g., “The clown fish is a popular kite design"), which is advantageous for A-OKVQA questions requiring external knowledge.
@@ 5 Related Work
Multi-hop Reasoning. Chain-of-thought (CoT) reasoning can be regarded as a form of multi-hop reasoning, as it involves constructing a sequence of reasoning to derive the answer. However, as indicated in Table 1, the traditional CoT method often results in erroneous information (e.g., visual hallucinations), leading to incorrect answers. Several recent works Dhuliawala et al. (2023); Jiang et al. (2023); Chen et al. (2023); Zelikman et al. (2022) have thus attempted to enhance the CoT reasoning capabilities. CoVe Dhuliawala et al. (2023) constructs a set of verification questions and uses them to verify the correctness of rationales. FLARE Jiang et al. (2023) iteratively identifies flawed rationales and corrects them by utilizing relevant documents from the knowledge base. CURE Chen et al. (2023) uses feedback from LLMs to tackle the hallucination during the generation process. STaR Zelikman et al. (2022) improves the model’s reasoning capability by supervising rationale generation based on the inference results of the train sets. In contrast, our II-MMR proposes a self-correction mechanism, which uses the initial prediction as the input context and autonomously corrects it if incorrect through the reasoning steps.
@@ 6 Conclusion
In this work, we propose II-MMR, a novel method to improve the reasoning capabilities of V&L models for VQA. Due to our constrained computational budget, we only conduct our studies on BLIP-2. We plan to extend our work to recent powerful V&L models such as LLaVA-1.5 Liu et al. (2023) or GPT-4V OpenAI2023 (2023). Besides, we conduct a small-scale human study (involving 3 annotators) to assess the quality of our rationales. The evaluation may be subjective among annotators due to the nature of the language. For instance, each annotator may have different opinions about the number of reasoning steps required for the same question. We intend to expand the scale of the human study to mitigate the subjectivity issue.
@@
  • [itemsep=5pt,topsep=1.5pt]
  • Appendix A: More details about dataset and training (cf. §4).
  • Appendix B: Qualitative results of incorrect prediction (cf. §4.3).
  • Appendix C: Fine-tuning details (cf. §4.4).
  • Appendix D: Augmenting questions with more reasoning complexity (cf. §4.5).
  • Appendix E: Additional Results (cf. §4).
@@ Appendix A More dataset and training details
Training. We mainly use BLIP-2 as the V&L model for our experiments. Concretely, we opt for the largest BLIP-2 model, which features the ViT-g/14 vision encoder from EVA-CLIP Fang et al. (2023), coupled with FlanT5XXL Chung et al. (2022), an encoder-decoder-based LLM. BLIP-2 incorporates a transformer-based bridge module that connects the vision encoder to the LLM. We focus on training the LLM and the bridge component while keeping ViT frozen. Our training configuration entails a batch size of 16, a learning rate of 1e-5, a beam size of 5, a maximum sequence length of 512, and an image resolution of 490 for ten epochs. We utilize eight A100 48GB GPUs for both training and inference.
@@ Appendix C Fine-tuning details
Appendix D Augmenting questions with more reasoning complexity
We see that the current VQA benchmarks lack multi-hop reasoning questions (cf. Table 2). We thus increase the reasoning complexity of their original questions and evaluate BLIP-2 on these new samples to better understand its reasoning capabilities. Concretely, we first collect questions that the model already answers correctly. Next, we extract keywords from those questions and use them as search queries to retrieve their relevant information from a large-scale textual knowledge base (i.e., Wikipedia). Leveraging this external knowledge, we ask an LLM to expand the current questions with more complex reasoning.
Appendix E Additional results

### v3

@@ Abstract
Visual Question Answering (VQA) often involves diverse reasoning scenarios across Vision and Language (V&L). Most prior VQA studies, however, have merely focused on assessing the model’s overall accuracy without evaluating it on different reasoning cases. Furthermore, some recent works observe that conventional Chain-of-Thought (CoT) prompting fails to generate effective reasoning for VQA, especially for complex scenarios requiring multi-hop reasoning. In this paper, we propose II-MMR, a novel idea to identify and improve multi-modal multi-hop reasoning in VQA. In specific, II-MMR takes a VQA question with an image and finds a reasoning path to reach its answer using two novel language promptings: (i) answer prediction-guided CoT prompt, or (ii) knowledge triplet-guided prompt. II-MMR then analyzes this path to identify different reasoning cases in current VQA benchmarks by estimating how many hops and what types (i.e., visual or beyond-visual) of reasoning are required to answer the question. On popular benchmarks including GQA and A-OKVQA, II-MMR observes that most of their VQA questions are easy to answer, simply demanding “single-hop” reasoning, whereas only a few questions require “multi-hop” reasoning. Moreover, while recent V&L models struggle with such complex multi-hop reasoning questions even using the traditional CoT method, II-MMR shows its effectiveness across all reasoning cases in both zero-shot and fine-tuning settings. [^1] https://github.com/heendung/II-MMR
@@
Jihyung Kil1thanks:  Work was partially done during the Amazon internship.11footnotemark: 1  Farideh Tavazoee2  Dongyeop Kang3  Joo-Kyung Kim2 1The Ohio State University  2Amazon AGI  3University of Minnesota  kil.5@osu.edu {fayt,jookyk}@amazon.com  dongyeop@umn.edu
@@ 1 Introduction
Effectiveness of II-MMR. II-MMR analyzes this reasoning path to identify different reasoning cases in current VQA benchmarks by measuring the number of reasoning steps and the types of reasoning, such as visual or beyond-visual (e.g., commonsense, knowledge base Schwenk et al. (2022)), required to answer the question (Figure 1). During our prompting process, the intermediate reasoning steps and the alignment of question keywords with visual objects determine the number and types of reasoning.
@@ 3 Experimental Setup
Training details. We follow the official configurations and implementations of BLIP-$2$https://github.com/salesforce/LAVIS/tree/main. Specifically, we select the largest BLIP-2 model, which its vision encoder is ViT-g Fang et al. (2023) and its LLM is FlanT5XXL Chung et al. (2024). During fine-tuning, the parameters of BLIP-2 are optimized with the language modeling loss on the downstream VQA training data with a batch size of 16 and a learning rate of 1e-5 for ten epochs. We utilize eight A100 48GB GPUs for both training and inference. See the appendix for more details.
@@ 4.1 Analysis of reasoning in VQA benchmarks
LLaVA-1.5+CoT 67.41 56.16 43.40 61.16 - - -
LLaVA-1.5+II-MMRApCoT 70.47 57.51 46.60 63.42 - - -
Table 6: Effectiveness of II-MMR over different reasoning cases in zero-shot VQA. Our II-MMR outperforms the traditional CoT on every reasoning case. Moreover, II-MMR notably improves over the baseline (BLIP-2) in all reasoning cases on both benchmarks, suggesting the benefit of our answer reasoning path in finding correct answers. Moreover, II-MMR shows its applicability to a recent VLM, LLaVA-1.5 Liu et al. (2023). Following Liu et al. (2023), we do not evaluate its zero-shot performance on A-OKVQA, as A-OKVQA was used during model training.
@@ 4.3 Benefit of II-MMR in zero-shot stage
Applicability of II-MMR. In addition to BLIP-2, we evaluate the effectiveness of our II-MMR on a more recent VLM, LLaVA-1.5 Liu et al. (2023). As shown in Table 6, II-MMR consistently improves over the traditional CoT in all reasoning cases, demonstrating its applicability to various VLMs with different architectural designs.
@@ 4.4 Benefit of II-MMR in fine-tuning stage
4.5 Expanding questions with more reasoning
As depicted in Table 2, the number of complex reasoning questions (i.e., 2-hop reasoning) is marginal in the GQA test-dev set, comprising only 3.73%. We thus conduct an ablation study: increasing the number of hops for each original question using a large-scale knowledge-base (e.g., Wikipedia) and evaluating the model performance on these newly expanded questions. We first extract keywords from the original question and use them as queries to retrieve their relevant information from Wikipedia. We then input this extra information (with the original question) into the LLM to increase the reasoning complexity of the question.
@@ 4.6 Qualitative Results
Figure 7 summarizes qualitative examples provided by our II-MMRApCoT. The generated rationale is highly relevant to the correct answer, leading the model to make the correct prediction. Moreover, our rationales entail knowledge beyond images, such as commonsense (e.g., “The clown fish is a popular kite design"), which is advantageous for A-OKVQA questions requiring external knowledge.
@@ 5 Related Work
Multi-hop Reasoning. Chain-of-thought (CoT) reasoning can be regarded as a form of multi-hop reasoning, as it involves constructing a sequence of reasoning to derive the answer. However, as indicated in Table 1, the traditional CoT method often results in erroneous information (e.g., visual hallucinations), leading to incorrect answers. Several recent works Dhuliawala et al. (2023); Jiang et al. (2023); Chen et al. (2024); Zelikman et al. (2022) have thus attempted to enhance the CoT reasoning capabilities. CoVe Dhuliawala et al. (2023) constructs a set of verification questions and uses them to verify the correctness of rationales. FLARE Jiang et al. (2023) iteratively identifies flawed rationales and corrects them by utilizing relevant documents from the knowledge base. CURE Chen et al. (2024) uses feedback from LLMs to tackle the hallucination during the generation process. Prophet Yu et al. (2023) is related to our work as it first obtains answer candidates and then includes them in the LLM prompt for the final prediction. However, Prophet relies on separate models for answer candidate selection (VQA) and final prediction (GPT-3 Brown et al. (2020)), which is less flexible than our single V&L model approach. Additionally, Prophet makes direct final predictions without generating rationales, thus lacking multi-hop reasoning capabilities. STaR Zelikman et al. (2022) improves the model’s reasoning capability by supervising rationale generation based on the inference results of the train sets. In contrast, our II-MMR proposes a self-correction mechanism, which uses the initial prediction as the input context and autonomously corrects it if incorrect through the reasoning steps.
@@ 6 Conclusion
In this work, we propose II-MMR, a novel method to improve the reasoning capabilities of V&L models for VQA. We conduct a small-scale human study (involving 3 annotators) to assess the quality of our rationales. The evaluation may be subjective among annotators due to the nature of the language. For instance, each annotator may have different opinions about the number of reasoning steps required for the same question. We plan to expand the scale of the human study to mitigate this issue.
@@ Appendix A More dataset and training details
Training. We mainly use BLIP-2 as the V&L model for our experiments. Concretely, we opt for the largest BLIP-2 model, which features the ViT-g/14 vision encoder from EVA-CLIP Fang et al. (2023), coupled with FlanT5-XXL Chung et al. (2024), an encoder-decoder-based LLM. BLIP-2 incorporates a transformer-based bridge module that connects the vision encoder to the LLM. We focus on training the LLM and the bridge component while keeping ViT frozen. Our training configuration entails a batch size of 16, a learning rate of 1e-5, a beam size of 5, a maximum sequence length of 512, and an image resolution of 490 for ten epochs. We utilize eight A100 48GB GPUs for both training and inference.
@@ Appendix C Fine-tuning details
Appendix D Additional results